OpenStudy (anonymous):

Kyle Estimates that his business in growing at a rate of 5% per year. HIs profits in 2005 were $67000. Estimate the profits in 2010.

OpenStudy (anonymous):

Is it along the same lines?

OpenStudy (johnweldon1993):

okay so we have exponential growth this time...meaning 1 + rate...so yeah basically the same thing

OpenStudy (johnweldon1993):

so you have the inital the rate the time all you need so the equation would look like?

OpenStudy (anonymous):

67000(1.05)^t?

OpenStudy (johnweldon1993):

correct...but what is t?

OpenStudy (anonymous):

Oh the answer- 85510.87

mathslover (mathslover):

Yep I got it...

OpenStudy (johnweldon1993):

there we go!

mathslover (mathslover):

67000 * (1.05)^5 u r right @zaynahf

OpenStudy (johnweldon1993):

yup...that would be correct @mathslover and @zaynahf

mathslover (mathslover):

$ 85110.864

OpenStudy (anonymous):

I have a slightly different formula: \[A= P (1+\frac{r }{ n })^{n \times t}\] where r is the rate, n is the time, P is the principle amount and n is the # of times compounded Do i need to insert the n value?

mathslover (mathslover):

n = 5?

OpenStudy (anonymous):

wait sorry, t is time

OpenStudy (johnweldon1993):

that would be if it was compounded at a different rate i believe...like if it was quarterly

OpenStudy (johnweldon1993):

i believe anyways

mathslover (mathslover):

Oh yeah @johnweldon1993 is right, compounded annually ...

OpenStudy (johnweldon1993):

since we only wanted annually....we didn't need to account for "n" like in that equation

mathslover (mathslover):

n = 1 ...

OpenStudy (johnweldon1993):

so imagine "N" was number of times per year the money got accounted for...in this case...since it is annually...n would = 1

mathslover (mathslover):

Agreed.^

OpenStudy (anonymous):

So if i had an investment of 4000 at 4.5% interest compounded quarterly for 5 years would it be: \[A= 4000(1+\frac{ .045 }{ 3 })^{3(5)}\]

OpenStudy (anonymous):

Im confused what i would put for n in that case

OpenStudy (johnweldon1993):

umm quarters in the year....i think it would be 4 quarters in the year....in 3 month intervals...so i think N would = 4? anyone confirm or deny?

mathslover (mathslover):

No : \[\large{A = 4000(1 + \frac{4.5}{5})^({4} \times 5)}\] When we talk about counding quarterly than that means n would be 4

mathslover (mathslover):

^ \(\large{4000 (1 + \frac{4.5}{4} )^{4\times 5}}\)

OpenStudy (anonymous):

so it wouldnt be 4 quarters times 5 years, making it 20? it would just be 4?

OpenStudy (johnweldon1993):

right....so 4000(1+(.045/4)) ^ (4 times 5)

mathslover (mathslover):

why 0.045 ?

OpenStudy (anonymous):

making it adecimal, sine its a percent

OpenStudy (anonymous):

since*

mathslover (mathslover):

Oh I never did that.. I usually use this formula : \[\large{A = P (1+\frac{r}{100})^{nt} }\] That's why I got confused :)

OpenStudy (johnweldon1993):

right....yeah i usually use that as well @mathslover however i stayed in context for this one

OpenStudy (johnweldon1993):

and it makes since too...if it WAS 4.5/4...that would be 1.125....that added to 1...would be 2.125....that a HUGE interest rate for a quarterly compound interest right?

mathslover (mathslover):

@zaynahf is ur doubt clear now?

mathslover (mathslover):

Yeah @johnweldon1993 :)

OpenStudy (johnweldon1993):

it would make since that the interest for quarterly would be less than the annual compound

OpenStudy (anonymous):

Oh okay i see... well thanks! :)

mathslover (mathslover):

No problem. Though I didn't help :D

OpenStudy (johnweldon1993):

You're welcome :)

OpenStudy (johnweldon1993):

and @mathslover yeah you did...you both added context and confirmed....that's as helpful as anything in math lol

OpenStudy (anonymous):

Okay so would the answer be 5003.002?

OpenStudy (johnweldon1993):

that's indeed what I get

OpenStudy (anonymous):

okay well thanks again

OpenStudy (johnweldon1993):

no problem! :)